THE SWITCHING/GAUSSIAN METHOD FOR POLARIZABLE CONTINUUM MODELS WITH APPLICA-TIONS TO AQUEOUS MOLECULAR DYNAMICS AND VIBRATIONAL SPECTRA

\underline{ADRIAN W. G. LANGE}, JOHN M. HERBERT, 100 W. 18th Ave, Columbus, OH 43210.

We describe the recently-developed Switching/Gaussian (SWIG) formalism for polarizable continuum models (PCMs) in electronic structure theory and QM/MM calculations. Existing PCM implementations are well-established, yet still exhibit Coulomb singularities and/or discontinuities in potential energy surfaces, which result from the finite-element discretization of the solute–continuum interface. Our revised implementation eliminates these problems. We illustrate the robustness this approach using ab initio molecular dynamics and vibrational spectra calculations, for which existing PCM implementations fail. Furthermore, we present the extension of SWIG to unprecedentedly large systems, making it a viable implicit solvation method even for molecular mechanics calculations of biological macromolecules.